OpenStudy (anonymous):

solve the differential equation: dy/dx + (y^3 + 4e^xy)/(2e^x+3y^2)

OpenStudy (anonymous):

can use bernoulli's method, linear, separable, homogenity test, etc.

OpenStudy (amistre64):

dy/dx + (y^3 +4e^xy)/(2e^x +3y^2) = 0 dy = -(y^3 +4e^xy)/(2e^x +3y^2) dx dy = -y(y^2 +4e^x)/(2e^x +3y^2) dx dy/-y(y^2 +4e^x) = dx/(2e^x +3y^2) Oww its giving me a headache :) Does that help any?

OpenStudy (anonymous):

yes but dx can only have functions of x with it and the same for y so is it not separable?

OpenStudy (amistre64):

It might be separable, but I cant see it yet :)

OpenStudy (anonymous):

the e^x is the problem

OpenStudy (anonymous):

i have an idea multiply all of them then take integral :)

OpenStudy (anonymous):

this is y^3 +4e^xy =u du/dy=2e^x+3y^2 may be this

OpenStudy (amistre64):

This ones outta my league for now....